OpenStudy (sette):

What is the simplifie form of : [2x/3(x+3)]/[7x/(x+3)(x-5)]. I know that inorder to divide radicals you flip the second denominator and numerator so the new equation would be : [2x/3(x+3)]*[(x+3)(x-5)/7x] However I do not know how to simplify the rest . Help ?

OpenStudy (bahrom7893):

cancel out the (x+3)s [2x/3]*[(x-5)/(7x)] cancel out the xs [2/3][(x-5)/7] multiply them together: 2*(x-5)/21

OpenStudy (bahrom7893):

Please click on Good Answer if I helped. You can also hover over my name and click on Become a fan ;)!

OpenStudy (mertsj):

\[\frac{2x}{3(x+3)}\div\frac{7x}{(x+3)(x-5)}=\frac{2x}{3(x+3)}\times\frac{(x+3)(x-5)}{7x}\]

OpenStudy (mertsj):

Now cancel x and x+3 and get: \[\frac{2(x-5)}{21}\]

OpenStudy (anonymous):

[2x/3(x+3)]/[7x/(x+3)(x-5)]. = [2x/3(x+3)] * (x+3)(x-5)/ 7x] = 2 ( x- 5) / 21